# AgentOccam: A Simple Yet Strong Baseline for LLM-Based Web Agents

### v1

@@ 4.2 Observation Space Alignment
Figure 3: The components of our web navigation agent’s prompt. It includes a general instruction outlining the task, the desired output and available actions, as well as online task information providing the current goal, the agent’s past interactions, and the latest observations. Notably, the sections on previous interactions and current observation use the most tokens. These can be attributed to two main factors: the length of the pages and the extent of history span, with current observation primarily depending on page length and past interactions on both page length and history span.
@@
Figure 4: To align the task’s observation space with the base model’s pre-training, we condense a single-page length by removing unnecessary texts that repetitively describe the web page’s functionality and layout (step 1), and by identifying page elements relevant to the task for the agent to remember (step 2). Additionally, we optimize the agent workflow memory through a planning tree, viewing each new plan as a separate goal and excluding past steps’ information dedicated to previous plans to enhance memory conciseness (step 3).
@@
The observation space of web agents consists of task objectives, instructions, previous interaction, the current web text descriptions or screenshots (see Figure 3 and Appendix D for our agent). Among them, previous interactions and current web content consumes the most number of tokens, which scales with the length of a single page and the length of history. This often results in a long context window, which not only increases the LLM inference cost but also poses challenges for LLM to extract related information accurately. Therefore, our main goal in refining the observation is to address these two aspects. Additionally, the alignment of observations is outlined in Figure 4.
@@
Simplifying Web Page Observations. The content on web pages are represented in HTML or accessibility tree format in most text-only web agents. These formats are designed towards front-end loading and rendering, containing numerous formatting tokens making them lengthy and repetitive, as illustrated in Figure 4 Step 1. Our goal is to optimize the representation to make it more readable to LLMs in one single page. Specifically, we merge function-descriptive web elements (e.g., StaticText [761] ‘My Account’) with interactive elements that share the same label (e.g., link [1312] ‘My Account’). We then convert table and list blocks to Markdown, eliminating repetitive structural tokens (e.g., columnheader, gridcell). Consequently, we achieve a more concise representation while keeping the same information.
@@
Replaying Observation History Selectively. Taking observation history as input is important for decision-making agents to act consistently for tasks requiring long horizons, given that the observation state only contains partial information about the environment’s state. For web tasks, it is also important to include both observation and action history as some key information may not be displayed on the current page. However, the observation history will also significantly scale up the context length and increase reasoning difficulty as well as inference cost. We address this issue by only selecting the most important and related information on previous web page, according to two rules based on the “pivotal” nodes (defined later) and the planning tree.
@@
First, we observe that only small amount of content on a web page is pertinent to a specific task among several steps and is worth to replay in future steps. For example, in tasks requiring the agent to find all reviews within three months, it is unnecessary to keep other reviews or some unrelated links like Contact Us on the page. Thus we employ a simple rule to identify this small amount of content by leveraging the tree structure of web data (e.g. accessibility tree). To do this, we first instruct the agent to pinpoint the crucial web elements denoted as “pivotal” nodes, at the same time when the agent generates an action. The agent is then programmed to include only the pivotal nodes’ ancestor nodes (indicating their global hierarchy and position), sibling nodes (providing immediate context), and descendant nodes (offering detailed characteristics) in the future observations as illustrated in Figure 4 Step 2. This effectively narrows down the volume of data and level of noise passed to future context of LLM inference.
@@
Second, we observe that not all previous steps’ observation needs to be noted during the inference of future step. Thus we can leverage the planning tree generated by the agent itself to keep the agent’s focus sharp. Specifically, when the agent initiates a branch action to develop a new plan, we treat this new plan as a separate goal. Steps taken for earlier plans and their observations will be dismissed in the current plan’s observation window, as depicted in Figure 4 step 3. This allows the agent to focus only on information dedicated to the current plan for a sub-task.
5 Experimental resuts and analysis
@@
We utilize WebArena (Zhou et al., 2023b), an interactive web simulator, as our benchmark. WebArena consists of fully functional websites from four common domains: e-commerce platforms (OneStopShop), social forums for idea and opinion exchange (Reddit), collaborative software development (GitLab), and content management for creation and management of online data (online store management). The platform additionally includes utility tools: a map, a calculator and a scratchpad, and Wikipedia to enable human-like task-solving. The benchmark consists of 812 tasks generated from 241 templates. A template here is a parametric form of a task intent, allowing for multiple instantiations with different keywords. Each task is accompanied by a evaluator/reward function that programmatically checks the correctness of the final information with respect to the desired ground truth information [^2] We identified and corrected errors in the original evaluators, with details discussed in Appendix B. Our approach outperforms the baseline methods with both original or corrected evaluators.. We use GPT-4-turbo-2024-04-09 (Achiam et al., 2023) to build our AgentOccam.
@@
We compare AgentOccam with the following prior and concurrent work: 1) WebArena agent: the Chain-of-Thought (CoT) prompted agent included in the WebArena benchmark (Zhou et al., 2023b). 2) SteP (Sodhi et al., 2024): a stack-based approach on top of 14 human-written atomic strategies tailored to solving WebArena. 3) WebPilot (Zhang et al., 2024): a multi-agent, multi-level MCTS based agent that reports state-of-the-art overall performance on WebArena. 4) Agent Workflow Memory (AWM) (Wang et al., 2024): a method automatically summarizing workflow from past experience. SteP has made their code and interaction trajectories public. Hence, we are able to fully replicate the agents from WebArena and SteP with GPT-4-turbo in the identical web environments as our methods, for a fair comparison. [^3] In our experiments, we note that all agents occasionally fails due to errors from the WebArena simulator, such as posting rate limits in Reddit or login expiration. In such cases, we restart the experiments. WebPilot and AWM, being concurrent works with this paper, have not yet provided source code or resulting trajectories, limiting our analysis of these works to just reporting the aggregated performance numbers included in their technical reports. Our analysis focuses on SteP as it is the most performant method prior to this work.
@@
Figure 5: Ablation study of AgentOccam’s action and observation space refinement. We incrementally add refinement components and evaluate their marginal performance gains.
@@
i) Removal of non-essential actions ($\downarrow$ Actions): Narrowing the action space can reduce the level of distraction for LLM policies and significantly improves performance across all tested websites as shown in Figure 5. By removing rarely used actions like tab_focus,go_forward, hover and press, the agent spends less steps wandering around and explores more efficiently using actions such as click and type. Table 3 shows it reduces hundreds of hover and goto actions while significantly increase the number of click and type.
ii) Disabling scrolling (Above + X Scrolling): We observe that LLM policies tend to use scroll up and down often when they do not know what to do (since these action are revertible). Consequently, it significantly delays the task execution and causes looping over in certain tasks. As a result, disabling the scrolling action and passing the entire page to agent proves advantageous, especially for GitLab and Reddit tasks. However, this strategy increases the number of observation tokens, which will be addressed by subsequent refinements.
@@
iv) Selective replay of web elements in one page (Above + History): In this experiment, we follow step 2 shown in Figure 4 to add a subset of elements from previous web pages as history. We observe that it allows the agent to avoid repetitive actions in tasks, significantly decreasing the steps needed for task completion as demonstrated in Table 5. However, this addition slightly hurts performance in tasks with dense single-page content or those requiring navigation across multiple pages, as shopping and Reddit tasks success rate drops by 3.2 and 6.0 points.
v) Planning via generation and selective replay of past pages (AgentOccam; Above + Planning): We introduce actions branch and prune to generate actions and exclude historical steps not in the current sub-plan from the current prompt context. This results in performance gains in tasks across nearly all websites, alongside a reduction in the required observation tokens. The actions branch and prune are both primarily used in correcting a failed strategy and trying an alternative path. For example, in the task of identifying the nearest national park to Boston (task id 265), the agent employs a branch action to adopt an alternative search strategy after a failed search attempt. In a GitLab task (task id 563), after multiple failed attempts using the Create project button, the agent opts for a prune action to explore other methods.
@@
The judge method is motivated by our observation of the high variation from the agent’s behavior. In some key steps, the agent has certain probability of generating the correct action but often failing to do so, making it hard for the agent to recover from later pages. For instance, when tasked with identifying the most suitable subreddit for posting (task template 6100), the AgentOccam agent tends to hastily choose less relevant subreddits and gets stuck there. To address this, we direct the AgentOccam to generate all possible suitable actions instead of one action at each step. These action candidates are then evaluated by another LLM (GPT-4-turbo as well) prompted to be play the role of a judge and select the best action. The prompts for the judge are included in Appendix D.
@@
By analyzing the trajectories of each method, we observe that the task-specific strategy like SteP can help when the strategy fits the task requirement. For example, in the task of "Draft an email to the shop owner via their contact us function for a coupon as {reason}" (task template 163), the AgentOccam + SteP and SteP agents excel by prompting the agent explicitly not to click the submit button after drafting, where AgentOccam fails to follow. However, for tasks outside the designed strategies, these hints can mislead the agent, leading to 2 points drop in overall success rate of AgentOccam + SteP compared to AgentOccam only. An example is task 639, where the agent, guided by SteP’s instruction "Under forums, you will see only a subset of subreddits. To get the full list of subreddits, you need to navigate to the Alphabetical option.", repetitively navigates away from the appropriate subreddit, and generates reasons for its action selection that "Clicking on the ’Alphabetical’ link will help us access a more comprehensive Reddit list.", demonstrating how hard-coded strategies can distract the agent and hurt generalizability.
The AgentOccam + Judge agent, combining the AgentOccam’s generated action list with the second opinion from a LLM judge increases its overall success rate by 2.6%, by completing tasks where it may well fail due to intermediate decision flaws. For example, in choosing the right subreddit for a post (task template 6100), the base AgentOccam might hastily pick from an initial list, whereas the AgentOccam + Judge agent conducts a thorough search using post keywords or explores the entire forum list before drafting the post. This approach minimizes errors due to rushed decisions, increasing the likelihood of successfully completing task series.
@@ 6 Conclusion
In this paper, we proposed a simple but efficient LLM-based web agent AgentOccam that refines its action and observation spaces to be more comprehensible for LLMs primarily trained on static text. Unlike other methods, AgentOccam stands out for its remarkably simple policy workflow, requiring no extra modules, LLM calls, or in-context examples. This simplicity does not compromise its performance; AgentOccam surpasses previous and contemporary approaches on WebArena by 9.8 (SteP) and 5.9 (WebPilot) absolute points, respectively. Our result emphasize the importance of maintaining a simple agent architecture for better generalizability, echoing the Occam’s razor principle. In summary, AgentOccam aims to lay a solid foundation and offer valuable insights for future web agent research and development.

### v2

@@ 4.2 Observation Space Alignment
Figure 3: The components of our web navigation agent’s prompt. It includes a general instruction outlining the task, the desired output and available actions, as well as online task information providing the current goal, the agent’s past interactions, and the latest observations. Notably, the sections on previous interactions and current observation use the most tokens, and can be attributed to two main factors: the length of the pages and the extent of history span.
@@
Figure 4: To align the web task’s observation space with the format that the base model perceives most effectively, we condense a single-page length by removing unnecessary texts that repetitively describe the web page’s functionality and layout (step 1), and by identifying page elements relevant to the task for the agent to remember (step 2). Additionally, we optimize the agent workflow memory through a planning tree, viewing each new plan as a separate goal and excluding past steps’ information dedicated to previous plans to enhance memory conciseness (step 3).
@@
The observation space of web agents consists of task objectives, instructions, previous interaction, and the current web text descriptions or screenshots (see Figure 3 and Appendix G for our agent). Among them, previous interactions and current web content consume the most number of tokens, which scales with the length of a single page and the length of history. This often results in a long context window, which not only increases the LLM inference cost but also poses challenges for LLM to extract related information accurately. Therefore, our main goal in refining the observation is to address these two aspects. The alignment of observations is outlined in Figure 4.
@@
Simplifying Web Page Observations. The content on web pages is represented in HTML or accessibility tree format in most text-only web agents. These formats are designed towards front-end loading and rendering, containing numerous formatting tokens making them lengthy and repetitive, as illustrated in Figure 4 Step 1. Our goal is to optimize the representation to make it more readable to LLMs on one single page. Specifically, we merge function-descriptive web elements (e.g., StaticText [761] ‘My Account’) with interactive elements that share the same label (e.g., link [1312] ‘My Account’). We then convert table and list blocks to Markdown, eliminating repetitive structural tokens (e.g., columnheader, gridcell). Consequently, we achieve a more concise representation while keeping the same information.
@@
Replaying Observation History Selectively. Taking observation history as input is important for decision-making agents to act consistently for tasks requiring long horizons, given that the observation state only contains partial information about the environment’s state. For web tasks, it is also important to include both observation and action history as some key information may not be displayed on the current page. However, the observation history will also significantly scale up the context length and increase reasoning difficulty as well as inference cost. We address this issue by only selecting the most important and related information on the previous web pages, according to two rules based on the “pivotal” nodes (defined later) and the planning tree. We provide detailed examples of how these two techniques are implemented in Appendix B.
@@
First, we observe that only a small amount of content on a web page is pertinent to a specific task among several steps and is worth replaying in future steps. For example, in tasks requiring the agent to find all reviews within three months, it is unnecessary to keep other reviews or some unrelated links like Contact Us on the page. Thus we employ a simple rule to identify this small amount of content by leveraging the tree structure of web data (e.g. accessibility tree). To do this, we first instruct the agent to pinpoint the crucial web elements denoted as “pivotal” nodes, at the same time when the agent generates an action. The agent is then programmed to include only the pivotal nodes’ ancestor nodes (indicating their global hierarchy and position), sibling nodes (providing immediate context), and descendant nodes (offering detailed characteristics) in the future observations as illustrated in Figure 4 Step 2. This effectively narrows down the volume of data and level of noise passed to the future context of LLM inference.
@@
Second, we observe that not all previous steps’ observation needs to be noted during the inference of future steps. Thus we can leverage the planning tree generated by the agent itself to keep the agent’s focus sharp. Specifically, when the agent initiates a branch action to develop a new plan, we treat this new plan as a separate goal. Steps taken for earlier plans and their observations will be dismissed in the current plan’s observation window, as depicted in Figure 4 step 3. This allows the agent to focus only on information dedicated to the current plan for a sub-task.
5 Experimental results and analysis
Here, we detail experiments on WebArena (Zhou et al., 2023b), a web simulator benchmark. Further experiments with WebVoyager (He et al., 2024), a web benchmark based on real-world websites, are included in Appendix C. We show AgentOccam’s base model generalizability in Appendix D.
@@ 5 Experimental results and analysis
We utilize WebArena, an interactive web simulator, as our benchmark. WebArena consists of fully functional websites from four common domains: e-commerce platforms (OneStopShop), social forums for idea and opinion exchange (Reddit), collaborative software development (GitLab), and content management for creation and management of online data (online store management). The platform additionally includes utility tools: a map, a calculator, a scratchpad, and Wikipedia to enable human-like task-solving. The benchmark consists of 812 tasks generated from 241 templates. A template here is a parametric form of a task intent, allowing for multiple instantiations with different keywords. Each task is accompanied by an evaluator that programmatically checks the correctness of the final information with respect to the desired ground truth information [^4] We identified and corrected errors in the original evaluators, with details discussed in Appendix E. Our approach outperforms the baseline methods with both the original and the corrected evaluators.. We use GPT-4-turbo-2024-04-09 (Achiam et al., 2023) to build our AgentOccam.
@@
We compare AgentOccam with the following prior and concurrent work: 1) WebArena agent: the Chain-of-Thought (CoT) prompted agent included in the WebArena benchmark (Zhou et al., 2023b). 2) SteP (Sodhi et al., 2024): a stack-based approach on top of 14 human-written atomic strategies tailored to solving WebArena. 3) WebPilot (Zhang et al., 2024): a multi-agent, multi-level MCTS based agent that reports state-of-the-art overall performance on WebArena. 4) Agent Workflow Memory (AWM) (Wang et al., 2024): a method automatically summarizing workflow from past experience. SteP has made its code and interaction trajectories public. Hence, we are able to fully replicate the agents from WebArena and SteP with GPT-4-turbo in the identical web environments as our methods, for a fair comparison. [^5] In our experiments, we note that all agents occasionally fail due to errors from the WebArena simulator, such as posting rate limits in Reddit or login expiration. In such cases, we restart the experiments. WebPilot and AWM, being concurrent works with this paper, have not yet provided source code or resulting trajectories, limiting our analysis of these works to just reporting the aggregated performance numbers included in their technical reports. Our analysis focuses on SteP as it was the most performant method prior to this work.
@@
Figure 5: Ablation study of AgentOccam’s action and observation space alignment, with details in Table 17. We incrementally add refinement components and assess marginal performance gains.
@@
i) Removal of non-essential actions ($\downarrow$ Actions): Narrowing the action space can reduce the level of distraction for LLM policies and significantly improves performance across all tested websites as shown in Figure 5. By removing rarely used actions like tab_focus,go_forward, hover and press, the agent spends fewer steps wandering around and explores more efficiently using actions such as click and type. Table 3 shows it reduces hundreds of hover and goto actions while significantly increasing the number of click and type.
ii) Disabling scrolling (Above + X Scrolling): We observe that LLM policies tend to use scroll up and down often when they do not know what to do (since these actions are revertible). Consequently, it significantly delays the task execution and causes looping in certain tasks. As a result, disabling the scrolling action and passing the entire page to the agent proves advantageous, especially for GitLab and Reddit tasks. However, this strategy increases the number of observation tokens, which will be addressed by subsequent refinements.
@@
iv) Selective replay of web elements in one page (Above + History): In this experiment, we follow step 2 shown in Figure 4 to add a subset of elements from previous web pages as history. We observe that it allows the agent to avoid repetitive actions in tasks, significantly decreasing the steps needed for task completion as demonstrated in Table 5. However, this addition slightly hurts performance in tasks with dense single-page content or those requiring navigation across multiple pages, as shopping and Reddit tasks success rate drops by 3.2 and 6.0 points, respectively.
v) Planning via generation and selective replay of past pages (AgentOccam; Above + Planning): We introduce actions branch and prune to allow the agent to autonomously generate plans and exclude historical steps not in the current sub-plan from the prompt context. This results in performance gains in tasks across nearly all websites, alongside a reduction in the required observation tokens. The actions branch and prune are both primarily used in correcting a failed strategy and trying an alternative path. For example, in the task of identifying the nearest national park to Boston (task id 265), the agent employs a branch action to adopt an alternative search strategy after a failed search attempt. In a GitLab task (task id 563), after multiple failed attempts using the Create project button, the agent opts for a prune action to explore other methods.
@@
The judge method is motivated by our observation of the high variation in the agent’s behavior. In some key steps, the agent has a certain probability of generating the correct action but often fails to do so, making it hard for the agent to recover from later pages. For instance, when tasked with identifying the most suitable subreddit for posting (task template 6100), the AgentOccam agent tends to hastily choose less relevant subreddits and gets stuck there. To address this, we direct the AgentOccam to generate all possible suitable actions instead of one action at each step. These action candidates are then evaluated by another LLM (GPT-4-turbo as well) prompted to play the role of a judge and select the best action. The prompts for the judge are included in Appendix G.
@@
By analyzing the trajectories of each method, we observe that task-specific strategies like those introduced in SteP can help when the strategy fits the task requirements. For example, in the task of "Draft an email to the shop owner via their contact us function for a coupon as {reason}" (task template 163), the AgentOccam + SteP and SteP agents excel by prompting the agent explicitly not to click the submit button after drafting, where AgentOccam fails to follow. However, for tasks outside the designed strategies, these hints can mislead the agent, leading to a 2-point drop in the overall success rate of AgentOccam + SteP compared to AgentOccam only. An example is task 639, where the agent, guided by SteP’s instruction "Under forums, you will see only a subset of subreddits. To get the full list of subreddits, you need to navigate to the Alphabetical option.", repetitively navigates away from the appropriate subreddit, and generates reasons for its action selection that "Clicking on the ’Alphabetical’ link will help us access a more comprehensive Reddit list.", demonstrating how hard-coded strategies can distract the agent and hurt generalizability.
The AgentOccam + Judge agent, combining the AgentOccam’s generated action list with the second opinion from an LLM judge increases its overall success rate by 2.6%, by completing tasks where it may well fail due to intermediate decision flaws. For example, in choosing the right subreddit for a post (task template 6100), the base AgentOccam might hastily pick from an initial list, whereas the AgentOccam + Judge agent conducts a thorough search using post keywords or explores the entire forum list before drafting the post. This approach minimizes errors due to rushed decisions, increasing the likelihood of successfully completing task series.
@@ 6 Conclusion
In this paper, we proposed a simple but effective LLM-based web agent AgentOccam that refines its action and observation spaces to be more comprehensible for LLMs primarily trained on static text. Unlike other methods, AgentOccam stands out for its remarkably simple policy workflow, requiring no extra modules, additional LLM calls, or in-context examples. This simplicity does not compromise its performance; AgentOccam surpasses previous and contemporary approaches on WebArena by 9.8 (SteP) and 5.9 (WebPilot) absolute points, respectively. Our results emphasize the importance of maintaining a simple agent architecture for better generalizability, echoing Occam’s razor principle. In summary, AgentOccam aims to lay a solid foundation and offer valuable insights for future web agent research and development.